# Scattering Amplitudes and Conservative Binary Dynamics at ${\cal O}(G^{5})$ without Self-Force Truncation

### v2

@@ Abstract
The IBP reduction is challenging due to the proliferation of irreducible numerators, doubled propagators, intricate topologies, and the large set of master integrals. We carry out the reductions with FIRE7 Smirnov and Zeng (2025), leveraging improved finite-field performance von Manteuffel and Schabinger (2015); Peraro (2016). In addition, we refine the choice of “seed integrals,” extending our previous strategy Bern et al. (2024b) and related developments Driesse et al. (2024); Guan et al. (2025). (see Kira 3 Lange et al. (2025) for a public implementation, Ref. Brunello:2025gpf for a recent application in the post-Newtonian expansion, and Refs. Larsen and Zhang (2016); Wu et al. (2023); Abreu et al. (2017, 2018) for analogous seeding schemes in earlier non-Laporta approaches.)
@@
We find that the coefficient of one classical 2SF eikonal sum, $\mathcal{E}_{2,15}$, associated with a Calabi–Yau threefold integral Brammer et al. (2025); Frellesvig et al. (2025); Duhr et al. (2025), vanishes to sufficiently high order in $\epsilon$ so that it does not contribute to the scattering angle. Similar cancellations have been observed at 5PM in maximal supergravity Bern et al. (2024b, 2025a), 4PN $\mathcal{O}(G^{5})$ Foffa et al. (2017) and at 5PM 1SF in Einstein gravity Driesse et al. (2024). This cancellation will be further discussed elsewhere Bern et al. . In contrast to the maximal supergravity case Bern et al. (2025a), the coefficient of the $\mathcal{E}_{2,8}$ eikonal sum, associated with a Heun function Joyce (1973); Joyce and Delves (2004); Ronveaux (1995); Ablinger et al. (2018), or equivalently symmetric-square K3, Picard–Fuchs equation Ruf (2023); Brammer et al. (2025), does not vanish.
@@
Terms through $\mathcal{O}(\bar{\sigma}^{22})$ are included in the ancillary file GR_potential_radial_action.m. The ancillary file also contains the exact analytic expressions for the 0SF and 1SF contributions. In these expressions, elliptic contributions cancel, as observed earlier in Ref. Driesse et al. (2024). The remaining functions are expressible in terms of manifestly real cyclotomic harmonic polylogarithms (CPL) Ablinger:2011te; Bern et al. (2024b) (c.f. also Ref. Ablinger:2021fnc for real representations of functions beyond CPLs). We note here the presence of terms proportional to $\pi^{2}$, which are absent from the analogous results in ${\cal N}=8$ supergravity Bern et al. (2025a). In our exact 1SF potential radial action, as well as in the 1SF conservative results of Ref. Driesse et al. (2024), such terms originate in (multiple) polylogarithmic functions. This is unlike the 4PM case, where $\pi^{2}$ terms tag only elliptic integrals. It will be interesting to determine the special functions that come with $\pi^{2}$ factors at 2SF order.
@@
Discussion— We include in Fig. 3 plots of the finite part of the radial action that illustrate some of their features. The left panel illustrates the relative importance of the different self-force components in the finite part of the radial action as a function of velocity. The right panel indicates that for $v\gtrsim.8$, higher-order terms and resummations will improve convergence. Once the radiation-mode contributions are incorporated, it will be essential to obtain accurate representations of the series solutions across the entire physical region. Several complementary strategies are available. As emphasized in Ref. Bern et al. (2025a), the choice of expansion variables plays a critical role in convergence. The conventional PN expansion in the velocity $v$ is not optimal, as it can introduce limitations from singularities outside the physical domain. A more suitable choice of expansion variable is $x$, related to the boost variable by $\sigma=(1+x^{2})/(2x)$, which leads to better convergence properties in the physical region $x\in[0,1]$. This strategy is incorporated in Fig. 3. See Ref. Ruf:Amplitudes23 for a discussion of spurious singularities of the solution to the differential equations.
@@
After obtaining the four-scalar integrand and performing the IBP reduction, the differential equations and boundary integrals completely fix the PM amplitude and scattering angle. We explicitly solved the differential equations for all master integrals as a series expansion about the static limit $\bar{\sigma}\to 0$. The resulting solution determines the classical amplitude (radial action) at $\mathcal{O}(G^{5})$ through 26PN order. Extending the series to substantially higher orders is, of course, doable using the differential equation Eq. (7) and might be useful once radiation contributions are included. It would be interesting to promote the series solution to an exact analytic result, with full control over generalized polylogarithms, elliptic integrals, and Calabi–Yau integrals.

### v3

@@ Abstract
The IBP reduction is challenging due to the proliferation of irreducible numerators, doubled propagators, intricate topologies, and the large set of master integrals. We carry out the reductions with FIRE7 Smirnov and Zeng (2025), leveraging improved finite-field performance von Manteuffel and Schabinger (2015); Peraro (2016). In addition, we refine the choice of “seed integrals,” extending our previous strategy Bern et al. (2024b) and related developments Driesse et al. (2024); Guan et al. (2025). (see Kira 3 Lange et al. (2025) for a public implementation, Ref. Brunello et al. (2025) for a recent application in the post-Newtonian expansion, and Refs. Larsen and Zhang (2016); Wu et al. (2023); Abreu et al. (2017, 2018) for analogous seeding schemes in earlier non-Laporta approaches.)
@@
We find that the coefficient of one classical 2SF eikonal sum, $\mathcal{E}_{2,15}$, associated with a Calabi–Yau threefold integral Frellesvig et al. (2024); Klemm et al. (2024); Brammer et al. (2025); Frellesvig et al. (2025); Duhr et al. (2025), vanishes to sufficiently high order in $\epsilon$ so that it does not contribute to the scattering angle. Similar cancellations have been observed at 5PM in maximal supergravity Bern et al. (2024b, 2025a), 4PN $\mathcal{O}(G^{5})$ Foffa et al. (2017) and at 5PM 1SF in Einstein gravity Driesse et al. (2024). This cancellation will be further discussed elsewhere Bern et al. . In contrast to the maximal supergravity case Bern et al. (2025a), the coefficient of the $\mathcal{E}_{2,8}$ eikonal sum, associated with a Heun function Joyce (1973); Joyce and Delves (2004); Ronveaux (1995); Ablinger et al. (2018), or equivalently symmetric-square K3, Picard–Fuchs equation Ruf (2023); Klemm et al. (2024); Duhr et al. (2025); Brammer et al. (2025), does not vanish.
@@
Terms through $\mathcal{O}(\bar{\sigma}^{22})$ are included in the ancillary file GR_potential_radial_action.m. The ancillary file also contains the exact analytic expressions for the 0SF and 1SF contributions. In these expressions, elliptic contributions cancel, as observed earlier in Ref. Driesse et al. (2024). The remaining functions are expressible in terms of manifestly real cyclotomic harmonic polylogarithms (CPL) Ablinger et al. (2011); Bern et al. (2024b) (c.f. also Ref. Ablinger et al. (2021) for real representations of functions beyond CPLs). We note here the presence of terms proportional to $\pi^{2}$, which are absent from the analogous results in ${\cal N}=8$ supergravity Bern et al. (2025a). In our exact 1SF potential radial action, as well as in the 1SF conservative results of Ref. Driesse et al. (2024), such terms originate in (multiple) polylogarithmic functions. This is unlike the 4PM case, where $\pi^{2}$ terms tag only elliptic integrals. It will be interesting to determine the special functions that come with $\pi^{2}$ factors at 2SF order.
@@
Discussion— We include in Fig. 3 plots of the finite part of the radial action that illustrate some of their features. The left panel illustrates the relative importance of the different self-force components in the finite part of the radial action as a function of velocity. The right panel indicates that for $v\gtrsim.8$, higher-order terms and resummations will improve convergence. Once the radiation-mode contributions are incorporated, it will be essential to obtain accurate representations of the series solutions across the entire physical region. Several complementary strategies are available. As emphasized in Ref. Bern et al. (2025a), the choice of expansion variables plays a critical role in convergence. The conventional PN expansion in the velocity $v$ is not optimal, as it can introduce limitations from singularities outside the physical domain. A more suitable choice of expansion variable is $x$, related to the boost variable by $\sigma=(1+x^{2})/(2x)$, which leads to better convergence properties in the physical region $x\in[0,1]$. This strategy is incorporated in Fig. 3. See Ref. Ruf for a discussion of spurious singularities of the solution to the differential equations.
@@
After obtaining the four-scalar integrand and performing the IBP reduction, the differential equations and boundary integrals completely fix the PM amplitude and scattering angle. We explicitly solved the differential equations for all master integrals as a series expansion about the static limit $\bar{\sigma}\to 0$. The resulting solution determines the classical amplitude (radial action) at $\mathcal{O}(G^{5})$ through 26PN order. Extending the series to substantially higher orders is, of course, doable using the differential equation Eq. (7) and might be useful once radiation contributions are included. It would be interesting to promote our series solution to an exact analytic result [^1] After this work was completed, such a solution was given in Ref. Driesse et al. (2026), which confirmed our potential region results. That paper also included conservative radiation effects., with full control over generalized polylogarithms, elliptic integrals, and Calabi–Yau integrals.